• Log in or sign up for an account here: https://www.sythe.org/login

Guide to staking 1m+ per hour GUARANTEED!

Status
Not open for further replies.

toykyo

Newcomer
Joined
Jan 9, 2011
Sythe Gold
0
Messages
16
Location
Boston, MA
I was at duel arena over the past week and I've developed a method for staking with guaranteed profit.

Now I have no idea if this method is already completely well known across the board, or if I just figured it out. PLEASE DO NOT CRITICIZE UNLESS YOU READ THE WHOLE THING BECAUSE IT ALL COMES TOGETHER, I'm not that great at guides, but I'm trying.

Anyway my theory works like this.

Requirements:

99 atk
99 str
99 def
99 constitution

preferably 100m starting cash, or 10m if you are willing to gain profit much slower.

Theory:
So in theory, boxing matches with maxed 99s are 50-50, sure someone could argue it may be 49-51, depending on the attack style, but for arguments sake lets use accurate, because hitting at least a little is the most important.

Method: (Step A is assuming that you lose many matches at first until you win)

Step A

Round 1: - Outcomes - (W) , (L)

Stake 1m

Now from this stake, it is 50-50. If you win, then you have a 1m profit. I will get to winning later. This part of the guide will be explaining losing.

So let's say you lose your first stake, you would be down 1m. If you win go to step B. I CANNOT STRESS THIS ENOUGH IF YOU WIN 1m DO NOT STAKE 2M the next stake!

Round 2: Possible outcomes - (L W) , (L L)

Stake 2m

Now for this stake, you stake 2m, as you stand to gain a profit of 1m if you win, and lose a net loss of 3m if you lose. Remember that it is still 50-50, but the chances of you losing two times in a row are 1/4, as 1/2 * 1/2 = 1/4. If you win, go to step B

Round 3: possible outcomes - (L L W) , (L L L)

Stake 4m

For this stake, if you stake 4m, you stand to gain a profit of 1m if you win, and lose a net loss of 7m if you lose. Chance of winning is still 50-50, but chance of you losing 3 in a row is 1 in 8. If someone wishes for me to explain the probabilities further I will do so in another post.
Again, if you win, go to step B

Round 4: possible outcomes - (L L L W) , (L L L L)

Stake 8m.

Now going in to this stake (if it has gotten this far and you are unlucky) you have lost 7m. By staking 8m, you stand to gain 1m profit by winning. The chance you would lose all 4 is 1 in 16, now I have to admit I like those odds.

Round 5+: (very unlikely)

Stake the total amount you have lost +1m so that if you win, you will always gain a profit of 1m.

STEP B - Win

Whenever you win, no matter if you had just staked 16m vs 16m and won, for a 1m profit, or won 1m profit on your 1st stake, ALWAYS start the process over. That way whenever you win you will make 1m profit, and losing will not matter, as it will always end in 1m profit.

RESULTS:

I did a mock test of this theory and came out with these results.

Started with 100m:

ROUND 1

stake 1m - lose
stake 2m - lose
stake 4m - win

profit 1m

ROUND 2

stake 1m - win

profit 1m

Round 3

stake 1m - lose
stake 2m - lose
stake 4m - lose
stake 8m - lose
stake 16m - win

profit 1m

Round 4

stake 1m - win

As you can see I made a profit of 4m. This whole process took me about 30 minutes. Now although it is possible to make much more money staking at a much faster rate, this method is GUARANTEED.

Now this is just the first guide I have ever written and I would love feedback, or criticism, or anything helpful haha. thanks for reading this far :)
 
Will probably do this tomorrow for an hour and give you some feed back.
 
you may want to add if you have enough money,like the way i do it

10m
20m
40m
80m
160m
320m. if you loose here just quit staking for a while/quit the game.

btw, just a quick add on. today i lost about 14 stakes in a row. all for over 10m,and i was not doing this method,i was just doing 20m-50m stakes blindly.people say odds are 50/50,but its not 2/4.then 1/4 and so on... its 50/50 EACH FIGHT. just because you lost the previous fight,doesnt give you a advantage in the next fight. my best advice is if you hit a loosing streak,take a break for the night,calm down, and try some new strategys.i cant publicly give out my strategy for winning,but it didnt take long to notice that it worked. i wish you all the best of luck.
 
If you have a set amount of cash, you will end up breaking even due to the statistics, this method only works with an unlimited supply of money.
 
Pokémon Trainer
Two Factor Authentication User
Tier 1 Prizebox
This is a common trick, but it's a good one :)
It's pretty much if you're panic and don't like losing gold, except after a few losses, you need to really go big, and you can eventually get cleaned by a glitcher etc.
 
You always stake 1m more than you have lost total, so if you win you stand to gain a net income of 1m. Notice I state that although each match is 50-50, that chance that you will lose each one is not 50-50. The possible outcomes for a set of 2 matches that are 50-50 are as follows

W - W
W - L
L - W
L - L

The chance you lose both times in a row is 1/4. That is what makes my theory correct, other than duels not being 50-50 even
 
Thanks for sharing. Just keep double stake ur lose. I bet most likely out of 10 u win at least 1 time.
 
You do realize that, technically, it is not 50/50 as to whether you win or lose. Plus, as sarahs said, (or what he meant) is that the events are not mutually exclusive.
 
This is called the Roulette method, and is pretty well know. Although it seems good in theory (say you have 127m and are doing the 1m profit), there is a major flaw to it.

With 127m you can do 7 stakes (each time doubling) before you get cleaned.

The odds are 1/128 that you lose 127m, and 127/128 that you win 1m. Basically in the long run you will find yourself at 0.
 
i disagree imagine this lets say you have 100m and you lose 1m then 3m and so on if you lose 6 times you are cleaned ok so the odds of this are 1 in 64 now that doesent sound too bad but this means that for every 1m you win theres a 1 in 64 chance youll be cleaned now lets do this for 10m you have 3 chances and your too far to climb out so thats 1 in 8 so for after 8m earned theres a 100% statistical chance ull be cleaned
 
Status
Not open for further replies.
Runescape Gamblingrunescape gambling
Back
Top